28(132^(39x - 20)) = 12481
A 2.755
B0.345
C1.675
D2.345
E 0.545
F 4.235

5e^(2.1x - 2) = 14
A 1.443
B2.543
C5.343
D2.434
E 1.343
F 0.443

Answer by lwsshak3(11628)

28(132^(39x - 20)) = 12481
(132^(39x - 20)) = 12481/28=445.75
(39x-20)=log(445.75)/log(132)
39x=20+log(445.75)/log(132)
x=0.545 ans: E
..
5e^(2.1x - 2) = 14
e^(2.1x - 2) = 14/5
(2.1x-2)lne=ln(14/5)
(2.1x-2)=ln(14/5)
2.1x=ln(14/5)+2
1.443 Ans A